# Logarithmic or algebraic: roughening of an active Kardar-Parisi-Zhang surface

### v3

@@ Appendix C Renormalisation group calculations
$S=-\int_{{\bf x},t}\hat{h}D\hat{h}+\int_{{\bf x},t}\hat{h}\bigg{\{}\partial_{t}h-\nu\nabla^{2}h-\frac{\lambda}{2}(\bm{\nabla}h)^{2}-\lambda_{1}\frac{\nabla_{i}\nabla_{j}}{\nabla^{2}}(\nabla_{i}h\nabla_{j}h)-\lambda_{2}\frac{\nabla_{i}\nabla_{j}}{\nabla^{2}}e_{jm}(\nabla_{i}h\nabla_{m}h)\bigg{\}}.$ (25)
In momentum space action functional can be split into two parts. The first one is the Gaussian part $S_{1}$ and other one is the anharmonic part, containing the local ($\lambda$ term) and nonlocal ($\lambda_{1}$ and $\lambda_{2}$ terms) anharmonic terms, $S_{2}$. Where,
@@
In $S_{2}$, we defined ${\bf k=k_{2}+k_{3}}$ and $\omega=\omega_{2}+\omega_{3}$. A factor of $1/k^{2}$ in the terms with coefficients $\lambda_{1}$ and $\lambda_{2}$ indicate their nonlocal nature in real space. This form of $S_{2}$ holds in 2D; at $d>2$, we must set $\lambda_{2}=0$, since the chiral term is 2D specific. As discussed in the main text, these nonlocal terms are contributions to the surface velocity that depend nonlocally on the local surface fluctuations.
C.1 Linear theory results
@@
Figure 3: Diagramatic representation of two point functions.
Fig. 3 shows diagramatic representation of the propagators.
C.2 Corrections to $D$
@@
Fig. 4$(a)$ has a symmetry factor $2$ and contributes to the correction of $D$ as shown below,
$\frac{2\lambda^{2}}{2!\times 4}\int_{{\bf q},\Omega}\frac{[\textbf{q}\cdot(\textbf{k}-\textbf{q})]^{2}\times 4D^{2}}{\bigl{(}\Omega^{2}+\nu^{2}q^{4}\bigl{)}\bigl{(}\Omega^{2}+\nu^{2}({\bf k-q})^{4}\bigl{)}}$ (27)
$=\frac{\lambda^{2}D^{2}}{4\nu^{3}}\int\frac{d^{2}q}{q^{2}}$ (28)
Fig. 4$(b)$ has a symmetry factor $2$ and contributes to the correction of $D$ as shown below,
$\frac{\lambda_{1}^{2}}{2!}\times 2\int_{{\bf q},\Omega}\frac{k_{i}k_{j}}{k^{2}}q_{i}q_{j}\frac{k_{m}k_{n}}{k^{2}}q_{m}q_{n}$
×$\quad\quad\quad\quad\quad\times\frac{4D^{2}}{(\Omega-i\nu q^{2})^{2}(\Omega+i\nu q^{2})^{2}}$ (29)
$=\frac{\lambda_{1}^{2}D^{2}}{\nu^{3}}\times\frac{3}{8}\int\frac{d^{2}q}{q^{2}}$ (30)
Fig. 4$(c)$ has a symmetry factor $2$ and contributes to the correction of $D$ as shown below,
$\frac{\lambda_{2}^{2}}{2!}\times 2\int_{{\bf q},\Omega}\frac{k_{i}k_{j}}{k^{2}}e_{jm}q_{i}q_{m}\frac{k_{s}k_{t}}{k^{2}}e_{tr}q_{s}q_{r}$
×$\quad\quad\quad\quad\quad\quad\times\frac{4D^{2}}{(\Omega-i\nu q^{2})^{2}(\Omega+i\nu q^{2})^{2}}$ (31)
$=\frac{\lambda_{2}^{2}D^{2}}{8\nu^{3}}\int\frac{d^{2}q}{q^{2}}$ (32)
Fig. 4$(d)$ has a symmetry factor $2$ and contributes to the correction of $D$ as shown below,
$\frac{\lambda\lambda_{1}}{2}\times 2\frac{k_{i}k_{j}}{k^{2}}\int_{{\bf q}}q_{i}q_{j}\times q^{2}\frac{4D^{2}}{4\nu^{3}q^{6}}$ (33)
$\frac{\lambda\lambda_{1}D^{2}}{2\nu^{3}}\int\frac{d^{2}q}{q^{2}}$ (34)
Adding all these contributions, we obtain the total corrections to $D$:
@@ C.3 Corrections to $\nu$
Fig. 5$(a)$ represents the one-loop contribution proportional to $\lambda^{2}$, which has a symmetry factor of $8$ and is given by
$-\frac{\lambda^{2}}{4}\times\frac{8}{2!}\int_{{\bf q},\Omega}({\bf q}\cdot{\bf(k-q)})({\bf q\cdot k})\times\frac{2D}{\Omega^{2}+\nu^{2}q^{4}}$
×$\quad\quad\quad\quad\quad\quad\quad\quad\quad\quad\quad\quad\quad\times\frac{1}{i\Omega+\nu({\bf k-q})^{2}}$
$=-2D\lambda^{2}\int_{{\bf q},\Omega}q_{i}(k-q)_{i}q_{j}k_{j}\times\frac{1}{\Omega^{2}+\nu^{2}q^{4}}$
×$\quad\quad\quad\quad\quad\quad\quad\quad\quad\quad\quad\quad\quad\times\frac{-i}{\Omega-i\nu({\bf k-q})^{2}}$
After doing the $\Omega$ integral above result becomes,
$-\frac{2D\lambda^{2}}{2\nu^{2}}\int_{\bf q}\frac{q_{i}(k-q)_{i}q_{j}k_{j}}{q^{2}(q^{2}+({\bf k-q})^{2})}$
Above result takes the following form after symmetrizing i.e. ${\bf q}\rightarrow({\bf q+\frac{k}{2}})$,
$-\frac{\lambda^{2}D}{\nu^{2}}\int_{\bf q}\frac{(q+\frac{k}{2})_{i}(\frac{k}{2}-q)_{i}(q+\frac{k}{2})_{j}k_{j}}{({\bf q+\frac{k}{2}})^{2}\bigl{\{}{\bf(q+\frac{k}{2})}^{2}+({\bf\frac{k}{2}-q)}^{2}\bigl{\}}}$
$=\frac{\lambda^{2}D}{4\nu^{2}}\int_{\bf q}\frac{q^{2}q_{j}k_{j}+q^{2}k^{2}}{{\bf(q+\frac{k}{2})}^{2}(q^{2}+\frac{k^{2}}{4})}.$
@@
This may be evaluated by using the well-known relation Yakhot and Orzag (1986)
@@
Fig. 5$(b)$ has symmetry factor $8$ and represents the one-loop contribution proportional to $\lambda_{1}^{2}$ to $\nu$. This reads
$-\frac{\lambda_{1}^{2}}{2!}\times 8\int_{{\bf q},\Omega}\frac{k_{i}k_{j}}{k^{2}}(q_{i}(k-q)_{j})\frac{q_{m}q_{n}}{q^{2}}k_{m}(k-q)_{n}$
×$\quad\times\frac{2D}{\Omega^{2}+\nu^{2}({\bf k-q})^{4}}\times\frac{1}{-i\Omega+\nu q^{2}}$
@@
$-\frac{8\lambda_{1}^{2}}{2!}\times\frac{2D}{2\nu^{2}}\int_{\bf q}\frac{k_{i}k_{j}}{k^{2}}q_{i}(k-q)_{j}\frac{q_{m}q_{n}}{q^{2}}k_{m}(k-q)_{n}$
×$\quad\quad\quad\quad\quad\quad\quad\quad\quad\times\frac{1}{({\bf k-q})^{2}\bigl{(}q^{2}+({\bf k-q})^{2}\bigl{)}}$
Again we shift ${\bf q}\rightarrow({\bf q+\frac{k}{2}})$ to obtain
$-\frac{4\lambda_{1}^{2}D}{\nu^{2}}\frac{k_{i}k_{j}k_{m}}{k^{2}}\int_{\bf q}\frac{(q+\frac{k}{2})_{i}(\frac{k}{2}-q)_{j}(q+\frac{k}{2})_{m}}{({\bf q+\frac{k}{2}})^{2}({\bf\frac{k}{2}-q})^{2}(2q^{2}+\frac{k^{2}}{2})}$
×$\quad\quad\quad\quad\quad\quad\quad\times\frac{(q+\frac{k}{2})_{n}(\frac{k}{2}-q)_{n}}{({\bf q+\frac{k}{2}})^{2}({\bf\frac{k}{2}-q})^{2}(2q^{2}+\frac{k^{2}}{2})}$
@@
First two terms in the numerator of the above integral vanish as, $k_{i}k_{j}$ is symmetric but $q_{m}(q_{i}\frac{k_{j}}{2}-q_{j}\frac{k_{i}}{2})$ is antisymmetric under the interchange of $i,j$. Third term of the integrand can be expanded for small $k$, which then takes the following form $-q_{i}q_{j}q_{m}({\bf k\cdot q}-{\bf k\cdot q})$ and vanishes identically. Only fourth term contributes and can be evaluated using the identity (36). Thus above integral reduces to
@@
Fig. 5$(c)$ has a symmetry factor $4$ and represents one of the contributions proportional to $\lambda_{1}\lambda$ where the external leg with $\hat{h}$ is from the vertex $\lambda_{1}$. This contribution is
-- $\frac{\lambda_{1}\lambda}{2}\times 4\times\frac{k_{i}k_{j}}{k^{2}}\int_{{\bf q},\Omega}q_{i}(k-q)_{j}({\bf q\cdot k})$
×$\quad\times\frac{2D}{\Omega^{2}+\nu^{2}q^{4}}\times\frac{1}{-i\Omega+\nu({\bf k-q})^{2}}$
@@
-- $2\lambda\lambda_{1}\frac{k_{i}k_{j}k_{m}}{k^{2}}\int_{\bf q}q_{i}(k-q)_{j}q_{m}\times\frac{2D}{2\nu^{2}q^{2}\bigl{(}q^{2}+({\bf k-q})^{2}\bigl{)}}$
After shifting ${\bf q}\rightarrow({\bf q+\frac{k}{2}})$ above result becomes,
$-\frac{2\lambda\lambda_{1}D}{\nu^{2}}\frac{k_{i}k_{j}k_{m}}{k^{2}}\int_{\bf q}\frac{(q+\frac{k}{2})_{i}(\frac{k}{2}-q)_{j}(q+\frac{k}{2})_{m}}{({\bf q+\frac{k}{2}})^{2}(2q^{2}+\frac{k^{2}}{2})}$
$=-\frac{2\lambda\lambda_{1}D}{\nu^{2}}\frac{k_{i}k_{j}k_{m}}{k^{2}}\int_{\bf q}\frac{q_{i}(q_{m}\frac{k_{j}}{2}-q_{j}\frac{k_{m}}{2})-\frac{k_{i}}{2}q_{j}q_{m}-q_{i}q_{j}q_{m}}{({\bf q+\frac{k}{2}})^{2}(2q^{2}+\frac{k^{2}}{2})}$
@@
The first integral in the above line can be calculated using identity (36). The second integral can be done by using the following identity Yakhot and Orzag (1986)
$k_{i}k_{j}k_{m}k_{n}\int d^{d}qf(q^{2})q_{i}q_{j}q_{m}q_{n}=k_{i}k_{j}k_{m}k_{n}$
×$\times\frac{[\delta_{ij}\delta_{mn}+\delta_{im}\delta_{jn}+\delta_{in}\delta_{jm}]}{d(d+2)}\int d^{d}qf(q^{2})q^{4}.$ (39)
@@
Fig. 5$(d)$ has symmetry factor $4$ and represents the contribution originating from vertices $\lambda_{1}$ $\lambda$ where external leg $\hat{h}$ is from vertex $\lambda$. The contribution is,
$-\frac{\lambda\lambda_{1}}{2}\times 4\int_{{\bf q},\Omega}{\bf q\cdot(k-q)}\frac{(k-q)_{i}(k-q)_{j}}{(k-q)^{2}}q_{i}k_{j}$
×$\quad\times\frac{2D}{\Omega^{2}+\nu^{2}q^{4}}\times\frac{1}{-i\Omega+\nu({\bf k-q})^{2}}$ (41)
@@
Fig. 5$(e)$ has a symmetry factor $2$ and represents the contribution proportional to $\lambda_{2}^{2}$. It is given by
$\frac{\lambda_{2}^{2}}{2!}\times 2\frac{k_{i}k_{j}}{k^{2}}e_{jm}e_{tn}\int_{{\bf q},\Omega}\Bigl{(}-q_{i}(k-q)_{m}-(k-q)_{i}q_{m}\Bigl{)}$
×$\quad\times\frac{(k-q)_{s}(k-q)_{t}}{({\bf k-q})^{2}}\Bigl{(}q_{s}k_{n}+q_{n}k_{s}\Bigl{)}$
×$\quad\times\frac{2D}{\Omega^{2}+\nu^{2}q^{4}}\times\frac{1}{i\Omega+\nu({\bf k-q})^{2}}$ (43)
@@
C.4 One-loop corrections to $\lambda,\lambda_{1}$ and $\lambda_{2}$

### v4

@@ III Renormalisation group calculations
[…]
Action functional $S$ can be split into two parts. The first one is the Gaussian part $S_{1}$ and other one is the anharmonic part, containing the local ($\lambda$ term) and nonlocal ($\lambda_{1}$ and $\lambda_{2}$ terms) anharmonic terms, $S_{2}$. Here,
@@
We have defined $\bf k=k_{2}+k_{3}$} and $\omega=\omega_{2}+\omega_{3}$. A factor of $1/k^{2}$ in the terms with coefficients $\lambda_{1}$ and $\lambda_{2}$ indicate their nonlocal nature in real space. This form of $S_{2}$ holds in 2D; at $d>2$, we must set $\lambda_{2}=0$, since the chiral term is 2D specific. As discussed in the main text, these nonlocal terms are contributions to the surface velocity that depend nonlocally on the local surface fluctuations.
III.1 Linear theory results
@@
Figure 3: Diagrammatic representation of two point functions.
Fig. 3 shows diagrammatic representation of the propagators.
III.2 Corrections to $D$
@@
Fig. 4(a)(a) has a symmetry factor $2$ and contributes to the correction of $D$ as shown below,
$\frac{2\lambda^{2}}{2!\times 4}\int_{{\bf q},\Omega}\frac{\bigl({\bf(q+\frac{k}{2})}\cdot{\bf(\frac{k}{2}-q)}\bigl)^{2}\times 4D^{2}}{\Bigl(\Omega^{2}+\nu^{2}({\bf q+\frac{k}{2}})^{4}\Bigl)\Bigl(\Omega^{2}+\nu^{2}({\bf\frac{k}{2}-q})^{4}\Bigl)}$ (27)
$=\frac{\lambda^{2}D^{2}}{4\nu^{3}}\int\frac{d^{2}q}{q^{2}},$ (28)
neglecting subleading contributions. Fig. 4(b)(b) has a symmetry factor $2$ and contributes to the correction of $D$ as shown below,
$-\frac{\lambda_{1}^{2}}{2!}\times 2\int_{{\bf q},\Omega}\frac{k_{i}k_{j}}{k^{2}}({\bf q+\frac{k}{2}})_{i}({\bf\frac{k}{2}-q})_{j}\frac{k_{m}k_{n}}{k^{2}}({\bf q+\frac{k}{2}})_{m}\times({\bf q-\frac{k}{2}})_{n}\times\frac{4D^{2}}{\Bigl(\Omega^{2}+\nu^{2}{\bf(q+\frac{k}{2})}^{4}\Bigl)\Bigl(\Omega^{2}+\nu^{2}{\bf(\frac{k}{2}-q)}^{4}\Bigl)}$ (29)
$=\frac{\lambda_{1}^{2}D^{2}}{\nu^{3}}\times\frac{3}{8}\int\frac{d^{2}q}{q^{2}},$ (30)
again neglecting subleading contributions. Fig. 4(c)(c) has a symmetry factor $2$ and contributes to the correction of $D$ as shown below,
$-\frac{\lambda_{2}^{2}}{2!}\times 2\int_{{\bf q},\Omega}\frac{k_{i}k_{j}}{k^{2}}e_{jm}({\bf q+\frac{k}{2}})_{i}({\bf\frac{k}{2}-q})_{m}\frac{k_{s}k_{t}}{k^{2}}e_{tr}({\bf q+\frac{k}{2}})_{s}({\bf q-\frac{k}{2}})_{r}\times\frac{2D}{\Bigl(\Omega^{2}+\nu^{2}{\bf(q+\frac{k}{2})}^{4}\Bigl)}\times\frac{2D}{\Bigl(\Omega^{2}+\nu^{2}{\bf(\frac{k}{2}-q)}^{4}\Bigl)}$ (31)
$=\frac{\lambda_{2}^{2}D^{2}}{8\nu^{3}}\int\frac{d^{2}q}{q^{2}},$ (32)
again neglecting subleading contributions. Fig. 4(d)(d) has a symmetry factor $2$ and contributes to the correction of $D$ as shown below,
$-\frac{\lambda\lambda_{1}}{2}\times 2\frac{k_{i}k_{j}}{k^{2}}\int_{{\bf q},\Omega}({\bf q+\frac{k}{2}})_{i}({\bf\frac{k}{2}-q})_{j}\Bigl({\bf(q+\frac{k}{2})}\cdot{\bf(q-\frac{k}{2})}\Bigl)\times\frac{4D^{2}}{\Bigl(\Omega^{2}+\nu^{2}{\bf(q+\frac{k}{2})}^{4}\Bigl)\Bigl(\Omega^{2}+\nu^{2}{\bf(\frac{k}{2}-q)}^{4}\Bigl)}$ (33)
$=\frac{\lambda\lambda_{1}D^{2}}{2\nu^{3}}\int\frac{d^{2}q}{q^{2}},$ (34)
neglecting subleading contributions. Adding all these contributions, we obtain the total corrections to $D$:
@@ III.3 Corrections to $\nu$
Fig. 5(a)(a) represents the one-loop contribution proportional to $\lambda^{2}$, which has a symmetry factor of $8$ and is given by
$-\frac{\lambda^{2}}{4}\times\frac{8}{2!}\int_{{\bf q},\Omega}\Bigl({\bf(q+\frac{k}{2})}\cdot{\bf(\frac{k}{2}-q)}\Bigl)\Bigl({\bf(q+\frac{k}{2})\cdot k}\Bigl)\times\frac{2D}{\Omega^{2}+\nu^{2}{\bf(q+\frac{k}{2})}^{4}}\times\frac{1}{i\Omega+\nu({\bf\frac{k}{2}-q})^{2}}$
After performing the $\Omega$ integral above result becomes,
$-\frac{\lambda^{2}D}{\nu^{2}}\int_{\bf q}\frac{({\bf q+\frac{k}{2}})_{i}({\bf\frac{k}{2}-q})_{i}({\bf q+\frac{k}{2}})_{j}k_{j}}{({\bf q+\frac{k}{2}})^{2}\bigl\{{\bf(q+\frac{k}{2})}^{2}+({\bf\frac{k}{2}-q)}^{2}\bigl\}}=\frac{\lambda^{2}D}{4\nu^{2}}\int_{\bf q}\frac{q^{2}q_{j}k_{j}+q^{2}k^{2}}{{\bf(q+\frac{k}{2})}^{2}(q^{2}+\frac{k^{2}}{4})}.$
@@
This may be evaluated by using the well-known relation Yakhot and Orszag (1986)
@@
Fig. 5(b)(b) has symmetry factor $8$ and represents the one-loop contribution proportional to $\lambda_{1}^{2}$ to $\nu$. This reads
$-\frac{\lambda_{1}^{2}}{2!}\times 8\int_{{\bf q},\Omega}\frac{k_{i}k_{j}}{k^{2}}\Bigl(({\bf q+\frac{k}{2}})_{i}({\bf\frac{k}{2}-q})_{j}\Bigl)\frac{({\bf q+\frac{k}{2}})_{m}({\bf q+\frac{k}{2}})_{n}}{\bf(q+\frac{k}{2})^{2}}\times k_{m}({\bf\frac{k}{2}-q})_{n}\times\frac{2D}{\Omega^{2}+\nu^{2}({\bf\frac{k}{2}-q})^{4}}\times\frac{1}{-i\Omega+\nu({\bf q+\frac{k}{2}})^{2}}$
@@
$-\frac{4\lambda_{1}^{2}D}{\nu^{2}}\frac{k_{i}k_{j}k_{m}}{k^{2}}\int_{\bf q}\frac{({\bf q+\frac{k}{2}})_{i}({\bf\frac{k}{2}-q})_{j}({\bf q+\frac{k}{2}})_{m}}{({\bf q+\frac{k}{2}})^{2}({\bf\frac{k}{2}-q})^{2}(2q^{2}+\frac{k^{2}}{2})}\times\frac{({\bf q+\frac{k}{2}})_{n}({\bf\frac{k}{2}-q})_{n}}{({\bf q+\frac{k}{2}})^{2}({\bf\frac{k}{2}-q})^{2}(2q^{2}+\frac{k^{2}}{2})}$
@@
First two terms in the numerator of the above integral vanish as, $k_{i}k_{j}$ is symmetric but $q_{m}(q_{i}\frac{k_{j}}{2}-q_{j}\frac{k_{i}}{2})$ is antisymmetric under the interchange of $i,j$. Third term of the integrand can be expanded for small $k$, which then takes the following form $q_{i}q_{j}q_{m}({\bf k\cdot q}-{\bf k\cdot q})$ and vanishes identically. Only the fourth term contributes and can be evaluated using the identity (36). Thus above integral reduces to
@@
Fig. 5(c)(c) has a symmetry factor $4$ and represents one of the contributions proportional to $\lambda_{1}\lambda$ where the external leg with $\hat{h}$ is from the vertex $\lambda_{1}$. This contribution is
$-\frac{\lambda_{1}\lambda}{2}\times 4\times\frac{k_{i}k_{j}}{k^{2}}\int_{{\bf q},\Omega}({\bf q+\frac{k}{2}})_{i}({\bf\frac{k}{2}-q})_{j}\Bigl({\bf(q+\frac{k}{2})\cdot k}\Bigl)\times\frac{2D}{\Omega^{2}+\nu^{2}({\bf q+\frac{k}{2}})^{4}}\times\frac{1}{i\Omega+\nu({\bf\frac{k}{2}-q})^{2}}$
@@
$-\frac{2\lambda\lambda_{1}D}{\nu^{2}}\frac{k_{i}k_{j}k_{m}}{k^{2}}\int_{\bf q}\frac{({\bf q+\frac{k}{2}})_{i}({\bf\frac{k}{2}-q})_{j}({\bf q+\frac{k}{2}})_{m}}{({\bf q+\frac{k}{2}})^{2}(2q^{2}+\frac{k^{2}}{2})}=-\frac{2\lambda\lambda_{1}D}{\nu^{2}}\frac{k_{i}k_{j}k_{m}}{k^{2}}\int_{\bf q}\frac{q_{i}(q_{m}\frac{k_{j}}{2}-q_{j}\frac{k_{m}}{2})-\frac{k_{i}}{2}q_{j}q_{m}-q_{i}q_{j}q_{m}}{({\bf q+\frac{k}{2}})^{2}(2q^{2}+\frac{k^{2}}{2})}$
@@
The first integral in the above line can be calculated using identity (36). The second integral can be done by using the following identity Yakhot and Orszag (1986)
$k_{i}k_{j}k_{m}k_{n}\int d^{d}qf(q^{2})q_{i}q_{j}q_{m}q_{n}=k_{i}k_{j}k_{m}k_{n}\times\frac{[\delta_{ij}\delta_{mn}+\delta_{im}\delta_{jn}+\delta_{in}\delta_{jm}]}{d(d+2)}\int d^{d}qf(q^{2})q^{4}.$ (39)
@@
Fig. 5(d)(d) has symmetry factor $4$ and represents the contribution originating from vertices $\lambda_{1}$ - $\lambda$ where external leg $\hat{h}$ is from vertex $\lambda$. The contribution is,
$-\frac{\lambda\lambda_{1}}{2}\times 4\int_{{\bf q},\Omega}\Bigl({\bf(q+\frac{k}{2})}\cdot{\bf(\frac{k}{2}-q)}\Bigl)\frac{({\bf\frac{k}{2}-q})_{i}({\bf\frac{k}{2}-q})_{j}}{{\bf(\frac{k}{2}-q)}^{2}}\times({\bf q+\frac{k}{2}})_{i}k_{j}\times\frac{2D}{\Omega^{2}+\nu^{2}({\bf q+\frac{k}{2}})^{4}}\times\frac{1}{i\Omega+\nu({\bf\frac{k}{2}-q})^{2}}$ (41)
@@
Fig. 5(e)(e) has a symmetry factor $2$ and represents the contribution proportional to $\lambda_{2}^{2}$. It is given by
$-\frac{\lambda_{2}^{2}k_{i}k_{j}}{k^{2}}\int_{{\bf q},\Omega}\Bigl(({\bf q+\frac{k}{2}})_{i}({\bf\frac{k}{2}-q})_{m}+({\bf\frac{k}{2}-q})_{i}({\bf q+\frac{k}{2}})_{m}\Bigl)\times e_{jm}e_{tn}\frac{({\bf\frac{k}{2}-q})_{s}({\bf\frac{k}{2}-q})_{t}}{({\bf\frac{k}{2}-q})^{2}}\Bigl(({\bf q+\frac{k}{2}})_{s}k_{n}+({\bf q+\frac{k}{2}})_{n}k_{s}\Bigl)$
$\times\frac{2D}{\Omega^{2}+\nu^{2}({\bf q+\frac{k}{2}})^{4}}\times\frac{1}{i\Omega+\nu({\bf\frac{k}{2}-q})^{2}}$ (43)
@@
III.4 One-loop corrections to $\lambda,\lambda_{1}$ and $\lambda_{2}$